OpenStudy (anonymous):

-1/2 + -7/8 = -4/5 -1 3/8 -1 3/11 -3/5

OpenStudy (anonymous):

-1/2 + -7/8 = -4/8 + -7/8 = -11/8

OpenStudy (paxpolaris):

\[-{11 \over8}= -1\frac38\]